how do you solve 6(t-5=2)t+5
Found 3 solutions by stanbon, richwmiller, georgiabh:
Answer by stanbon(75887)

solve 6(t-5=2)t+5
Your parentheses are misplaced.
Cheers,
Stan H.

Answer by richwmiller(17219)

You have the ) in the wrong place for sure.
probably should be
6*(t-5)=(2t+5)
6t-30=2t+5
4t=25
t=25/4 =6.25
or 6*(t-5)=2(t+5) more likely
6t-30=2t+10
4t=40
t=10

Answer by georgiabh(4)

Your brackets seem to be in the wrong place.
Assuming that the question should be 6(t-5) = 2(t+5)
We want to find the value of t. To do this you need to get t on its' own on one side of the equation.
First you need to multiply out each side of the equation and lose the brackets.
6t - 30 = 2t + 10
Then you need to get all of the t's on one side and the numbers on the other.
Remember that what ever you do on one side you must do on the other
6t - 2t - 30 = 2t - 2t + 10
4t - 30 = 10
4t - 30 + 30 = 10 + 30
4t = 40
Now we have the t's on one side and the number on the other we can find out what t is.
To find out the value of t we divide each side by 4
4t/4 = 40/4
t=10
If you are ever worried about whether or not you have found the right answer, simply put it back into the original equation so in this case,
6(10-5) = 2(10+5)
6x5=30=2x15
Tip: If you have extra time in an exam always go back and do this just to make sure